PROCEEDINGS

OF THE INTERNATIONAL SEMINAR

ON INTEGRABLE SYSTEMS Dedicated to the memory of M. V. SAVELIEV

Held at the Max–Planck–Institut für Mathematik, Bonn, February 22, 1999 M. V. SAVELIEV 17.X.1945–20.IX.1998

Contents

Introduction P. Sorba. Deformed Virasoro and W–algebras from elliptic algebras. D. I. Olive. In memory of Misha Saveliev: Lie algebras, integrability and particle

physics.

J.–L. Gervais. Lax equations in ten–dimensional supersymmetric classical Yang– Mills theories.

Ch. Devchand, J. Schiff. Supersymmetric integrable systems from geodesic flows on superconformal groups

R. M. Kashaev. The pentagon equation and mapping class groups of punctured surfaces.

D. Leites, A. Sergeev. Orthogonal polynomials of discrete variable and Lie algebras of complex size matrices.

A. M. Levin, M. A. Olshanetsky. Non-autonomous Hamiltonian systems related to higher Hitchin integrals.

A. N. Leznov. New approach to representation theory of semisimple Lie algebras and quantum algebras.

V. S. Matveev, P. J. Topalov. Geodesic equivalence of metrics as a particular case of integrability of geodesic flows.

L. O’Raifeartaigh, J. M. Pawlowski and V. V. Sreedhar. On the duality of quantum Liouville Field Theory.

A.M. Perelomov. A remark on matrix Airy function. B. Shoikhet, A. Vershik. Lie algebras with continual root system and Cartan

subalgebra C[h].